Can you help me solve an multi-step equation
-219+10x=57-2x

Answer by Boreal(15235)

Put the x s on one side and the constants on the other.
-219+10x=57-2x
Add 2x to both sides
-219+12x=57
Add 219 to both sides
12 x = 276
divide by 12 NOT by 12x
x=276/12 =23
Check
-219 +(23*10)= 11
52-2(23) =11
Checks.